The Following are the height in centimeters and weights in kilogram of 5 teachers in a certain schools.Determine they relationship between the height (cm) and weight (kg)


Teacher A B C D E


HEIGHT (cm) X 163,160,168,159,170


Weight (kg) Y 52, 50, 64, 51, 69

1
Expert's answer
2022-06-13T18:03:04-0400

In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 163 & 52 & 8476 & 26559 & 2704 \\
 \hdashline
 & 160 & 50 & 8000 & 25600 & 2500 \\
 \hdashline
 & 168 & 64 & 10752 & 28224 & 4096 \\
 \hdashline
 & 159 & 51 & 8109 & 25281 & 2601 \\
 \hdashline
 & 170 & 69 & 11730 & 28900 & 4761 \\
 \hdashline
Sum= & 820 & 286 & 47067 & 134574 & 16662 \\
 \hdashline
\end{array}"




"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{820}{5}=164"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{286}{5}=57.2"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=134574-\dfrac{820^2}{5}=94"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=16662-\dfrac{(286)^2}{5}=302.8"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=47067-\dfrac{820(286)}{5}=163"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{163}{\sqrt{94(302.8)}}"




"=0.9662"


Strong positive correlation


"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{163}{94}=1.7340"




"n=\bar{Y}-m\bar{X}=57.2-\dfrac{163}{94}(164)=-227.1830"


The regression equation is:


"y=-227.1830+1.7340x"